# Locally-APN Binomials with Low Boomerang Uniformity in Odd Characteristic

### v1

@@ 4.2 Boomerang Spectrum of $F_{2}$ when $p=3$
In the above subsection, we showed that the boomerang uniformity of $F_{r}$ is at most 2 for several values of $r$, especially when $r=\frac{3^{k}+1}{2}$, that is, generalized Coulter-Matthews exponent. Next, we consider the special case $k=1$ and prove that the boomerang uniformity of $F_{r}$ is actually $1$ in this case. When $k=1$, we have $r=2$, and the boomerang uniformity of $F_{2}$ was studied in [14]. Theorem $8$ of [14] proved that the boomerang uniformity of $F_{2}$ equals $2$ for $q\geq 9613^{2}$, and conjectured that this holds for $q>307$. However, we demonstrate that this result does not hold when $p=3$ ; specifically, the boomerang uniformity is $1$ in this case. Moreover, we also provide the full boomerang spectrum of $F_{2}$.
@@
$\nu_{1}=\displaystyle\frac{1}{8}(2A-4)=\displaystyle\frac{1}{4}(A-2),$ (30)
@@
Let $y=b^{\frac{q+1}{4}}$. Then, applying Lemma 2.4 and Lemma 2.5,
@@
$=q-1-\displaystyle\sum_{y\in\mathbb{F}_{q}}\left(\chi(y(y-1))+\chi(y+1)+\chi(y(y^{2}-1))\right)\chi\left((y+1)^{\frac{q+1}{4}}-1\right).$
@@
$A$ $=q-1-2\displaystyle\sum_{\chi(t)=1}\chi((t^{2}-1)(t^{2}-2)(t-1))=q+1-2\displaystyle\sum_{\chi(t)=1}\chi((t+1)(t^{2}-2))$
@@
Let $z=\frac{t(t+1)}{t^{2}+1}$. Then, $z=1$ if and only if $t=1$. If $t=1$, then $(z-1)t^{2}-t+z=0$. The discriminant of this quadratic equation on $t$ can be computed as $(-1)^{2}-z(z-1)=1+z-z^{2}$. Hence, applying z=s1z=-s-1, we obtain
@@ 5 Conclusion
In this paper, we show that if $\gcd(q-1,r)\mid 2$ and $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$, then $F_{r}$ is locally-APN with boomerang uniformity at most $2$. Moreover, it holds that $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$ for the exponents $r$ listed in Table 3. We further investigate the differential spectra of $F_{3}$ and $F_{\frac{2q-1}{3}}$. In addition, by analyzing the boomerang spectrum of $F_{2}$ when $p=3$, we prove that its boomerang uniformity is equal to $1$, which contradicts the result of [14], where it was claimed that the boomerang uniformity of $F_{2}$ equals $2$ for sufficiently large $q$.
@@
Acknowledgements : This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2021R1C1C2003888). Soonhak Kwon was supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education (No. RS-2019-NR040081).

### v2

@@ 4.2 Boomerang Spectrum of $F_{2}$ when $p=3$
In the above subsection, we showed that the boomerang uniformity of $F_{r}$ is at most 2 for several values of $r$, especially when $r=\frac{3^{k}+1}{2}$, that is, generalized Coulter-Matthews exponent. Next, we consider the special case $k=1$, that is $r=2$, and prove that the boomerang uniformity of $F_{r}$ is actually $1$ in this case. Moreover, we also provide the full boomerang spectrum of $F_{2}$.
Remark 4.7.
The boomerang uniformity of $F_{2}$ was studied in [14]. Theorem 8 of [14] proved that the boomerang uniformity of $F_{2}$ equals $2$ for $q\geq 9613^{2}$, and conjectured that this holds for $q>307$. We remark that the conclusion $\beta_{F_{2}}=2$ in [14] does not hold when $p=3$ and $n$ is odd. Indeed, $\beta_{F_{2}}=1$ consistently holds in this case, as established in our Theorem 4.9. The issue is that, in characteristic $3$, the identity 1=21=-2 makes certain cases considered in [14] incompatible, although they are treated as independent (e.g., conditions arising in Corollary 4). This phenomenon is specific to characteristic $3$, where 1=21=-2, and does not occur for odd primes $p>3$.
@@
$\nu_{1}=\displaystyle\frac{1}{4}\displaystyle\sum_{b\in\mathbb{F}_{q}\setminus\mathbb{F}_{3}}\left(1+\chi(b)\chi\!\left(b^{\frac{q+1}{4}}-1\right)\right)\left(1+\chi\!\left(b^{\frac{q+1}{4}}+1\right)\right)\left(1-\chi\left(\left(b^{\frac{q+1}{4}}+1\right)^{\frac{q+1}{4}}-1\right)\right)=\displaystyle\frac{1}{4}(A-2),$ (29)
@@
Let $y=b^{\frac{q+1}{4}}$. Then, since $\frac{q+1}{4}$ is odd, applying Lemma 2.4 and Lemma 2.5,
@@
$=q-1-\displaystyle\sum_{y\in\mathbb{F}_{q}}\left(\chi(y(y-1))+\chi(y+1)+\chi(y(y^{2}-1))\right)\chi\left((y+1)^{\frac{q+1}{4}}-1\right)$
$=q-2-\displaystyle\sum_{y\in\mathbb{F}_{q}\setminus\{-1\}}\left(\chi(y(y-1))+\chi(y+1)+\chi(y(y^{2}-1))\right)\chi\left((y+1)^{\frac{q+1}{4}}-1\right).$
@@
$A$ $=q-1-2\displaystyle\sum_{\chi(t)=1}\chi((t^{2}-1)(t^{2}-2)(t-1))=q-1-2\displaystyle\sum_{\begin{subarray}{c}\chi(t)=1\\ t\neq 1\end{subarray}}\chi((t+1)(t^{2}-2)(t-1)^{2})$
$=q-1-2\displaystyle\sum_{\begin{subarray}{c}\chi(t)=1\\ t\neq 1\end{subarray}}\chi((t+1)(t^{2}-2))=q+1-2\displaystyle\sum_{\chi(t)=1}\chi((t+1)(t^{2}-2))$
@@
Let $z=\frac{t(t+1)}{t^{2}+1}$. Then, $z=1$ if and only if $t=1$. For $t\neq 1$, then $(z-1)t^{2}-t+z=0$. The discriminant of this quadratic equation on $t$ can be computed as $(-1)^{2}-z(z-1)=1+z-z^{2}$. Hence, applying z=s1z=-s-1, we obtain
@@ 5 Conclusion
In this paper, we show that if $\gcd(q-1,r)\mid 2$ and $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$, then $F_{r}$ is locally-APN with boomerang uniformity at most $2$. Moreover, it holds that $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$ for the exponents $r$ listed in Table 3. We further investigate the differential spectra of $F_{3}$ and $F_{\frac{2q-1}{3}}$. In addition, by analyzing the boomerang spectrum of $F_{2}$ when $p=3$, we prove that its boomerang uniformity is equal to $1$, which shows that the claimed conclusion in [14] does not hold in characteristic $3$, where it was claimed that the boomerang uniformity of $F_{2}$ equals $2$ for sufficiently large $q$.
@@
Acknowledgements: The authors would like to thank the anonymous reviewers for their kind comments and valuable suggestions. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2021-NR061794). Soonhak Kwon was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (No. RS-2019-NR040081).